Exactly 1.387 g of compound Z was combusted to form 3.499 g CO₂ mo sand 0.614 g H₂O. Compound Z contains only carbon, hydrogen, and oxygen. Determine the compound's molecular formula if the molar mass of compound Z is 122.118 g/mole.

1
Expert's answer
2022-04-22T04:09:03-0400

СxHyOz + O2 → xCO2 + x/yH2O


n(CO2) = m(CO2)/Mr(CO2) = 3.499 g / 44 g/mol = 0.0795 moll

It containd 0.0795 moll of C

m(C) = n(C)*Mr(C) = 0.0795 moll * 12 g/moll = 0.954 g


n( H2O) = m(H2O)/Mr(H2O) = 0.614 g / 18 g/mol = 0.0341 moll

It containd 2*0.0795 moll of H = 0.0682 moll

m(H) = n(H)*Mr(H) = 0.0682 moll * 1 g/moll = 0.0682 g


m(O) = m(СxHyO) - m(C) - m(H) = 1.387 g - 0.954 g - 0.0682 g = 0.3648 g

n(O) = m(O)/Mr(O) = 0.3648 g / 16 g/mol = 0.0228 moll


ratio n(C):n(H):n(O) = 0.0795 : 0.0682 : 0.0228 = 3.5:3:1 = 7:6:2


Molecular formula is С7H6O2